OpenStudy (anonymous):

Given that F(x) = x 2 + 2, evaluate F(1) + F(5).

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

yo

OpenStudy (nincompoop):

replace f with 1 and 5 add them

OpenStudy (anonymous):

Okay look here \[F(x) = x ^2 + 2\] You want to find F(1) + F(5), so --------------- F(1) means \[F(1)=1^2+2=3\] --------------- F(5) means \[F(5)=5^2+2=27\] -------------- So F(1)+F(5) means 3+27= 30 Got it?